# Offline-online approximation of multiscale eigenvalue problems with random defects

### v1

@@ 1 Introduction
In this work, we aim to obtain accurate and efficient approximations of the solutions to the stochastic PDE for many different samples. However, in a multiscale setting, standard finite element methods are, in general, not feasible due to the required fine discretization. Therefore, computational multiscale methods are employed, where we focus on the Localized Orthogonal Decomposition (LOD) method [MP21, HM14, EHMP19, MP14] in this work. The key idea is to construct problem-adapted basis functions incorporating information about the multiscale nature of the problem, so that a low-dimensional ansatz space yields sufficiently accurate solutions. However, solving the problem for many different samples remains quite expensive, as the problem-adapted basis functions must be re-computed for each realization of the (perturbed) coefficient.
The combined challenges of randomness and multiscale features have been tackled in several studies, often in the MC context and based on different multiscale methods. [Owh17] constructs so-called gamblets – hierarchical basis functions that enable sparse and orthogonal multiresolution decomposition. [ABS13] considers multilevel MC methods for problems with several, separated scales. In [BH17], the authors present a finite element-based hierarchical MC method for problems with periodic multiscale coefficients generated from an ergodic dynamical system acting on a probability space. [ZCH15] combines multiscale model reduction with data-driven approaches to obtain a reduced dimensional system in both physical and stochastic aspects. In [LBLT14], the authors propose a Multiscale Finite Element Method with a deterministic multiscale basis by taking into account the perturbative nature of the random coefficients, which have a similar structure as our model below. Very recently, [EKM24] suggests a combination of the LOD with neural networks for an efficient evaluation of the coefficient-to-basis map for general parametrized multiscale problems, that might be applicable in the random case as well. All the aforementioned works primarily consider the source problem. With regard to the eigenvalue problem, [XZO19] combines operator-adapted wavelets (gamblets) with hierarchical subspace correction to efficiently block-diagonalize the operator and refine solutions across scales. The algorithm computes eigenpairs through coarse-to-fine corrections in a multigrid fashion. [AP19] uses a preconditioned inverse iteration with an optimal multigrid solver to efficiently compute approximations of the localized low-energy eigenstates to a linear random Schrödinger operator. Note that we cannot expect such localized eigenstates in our setting.
@@
$-\nabla\cdot(A(x,\omega)\nabla u(x,\omega))=\lambda(\omega)u(x,\omega)\qquad\text{with}\qquad A(x,\omega)=A_{\varepsilon}(x)+b_{p}(x;\omega)B_{\varepsilon}(x).$ (1)
Detailed assumptions are postponed to the next Section 2.1. The random defects model for $A$ is the same as in [MV22] and originally considered in [ALB11, ALB12]. Figure 1 represents two example coefficients of $A$ generated by the background coefficient $A_{\varepsilon}$ arranged in $\varepsilon$-sized light-green blocks, and the inclusion coefficient $B_{\varepsilon}$ depicted by dark-green blocks. The random inclusion (Figure 1(a)) and erasure (Figure 1(b)) of a dark-green block represent a defect.
@@ 2.1 Model problem
In this section, we describe the problem for a single sample $\omega$. For simplicity, we thus omit the random variable $\omega$ in the following. Recall, however, that the final goal is to find an effective way of solving this problem for many samples. We assume the domain $\Omega\subset\mathbb{R}^{d}$ to be a unit hypercube, i.e., $\Omega=[0,1]^{d}$ with $d\in\{1,2,3\}$ where we impose periodic boundary conditions. When we later discretize our problem, we always use square meshes.
The weak form of the elliptic eigenvalue problem is given by: Find $(\lambda,u)\in\mathbb{R}\times V$ such that
@@
The functional space $V$ is the conforming finite element space given by
$V:=V_{h}\subseteq\mathcal{H}^{1}_{\#}(\Omega)=\{v_{1}\in L^{2}(\Omega):\;\nabla v_{1}\in L^{2}(\Omega),v_{1}\text{ is periodic on }\partial\Omega\},$
where we assume the mesh size $h$ to be small enough that we treat the solution $u\in V_{h}$ as the reference solution, i.e., the error towards the exact solution in #$\mathcal{H}^{1}_{\#}(\Omega)$ is negligible. The energy norm and the $L^{2}$-norm are as usually defined by $a(\cdot,\cdot)=:{\norm{\cdot}^{2}_{\mathcal{A}}}={\norm{A^{1/2}\nabla\cdot}^{2}_{L^{2}(\Omega)}}$ and $m(\cdot,\cdot)=:\norm{\cdot}^{2}_{L^{2}(\Omega)}$ respectively. The realization $A\in L^{\infty}(\Omega,\mathbb{R})$ is a sample of a randomly perturbed coefficient which is uniformly bounded and elliptic
@@
Furthermore, the underlying coefficients $A_{\varepsilon}$ and $B_{\varepsilon}$ of $A(x)$ are taken to be spectrally bounded as
@@
$0<\alpha\leq\text{ess}\inf\limits_{x\in\Omega}(A_{\text{per}}(x)+B_{\text{per}}(x)),\;\;\;\;\infty>\beta\geq\text{ess}\sup\limits_{x\in\Omega}(A_{\text{per}}(x)+B_{\text{per}}(x))(x)$
@@
Here, the function $\chi$ is the characteristic function of each defect with $Q\subseteq[0,1]^{d}$ and $I:=\{\kappa\in\mathbb{Z}^{d}\ |\varepsilon(\kappa+Q)\subset\Omega\}$. The independent random variables $\hat{b}^{i}_{p}$ are independently Bernoulli distributed with probability $p$. Precisely, $\hat{b}^{i}_{p}=0$ with probability $1-p$ and $\hat{b}^{i}_{p}=1$ with probability $p$.
@@
For conciseness, we use the notation $a\lesssim b$ to denote the relationship $a\leq Cb$ where the generic constant $C$ is independent of the discretization parameters $H,h,k$ as well as $\varepsilon$ throughout the rest of the paper.
@@ 2.2 The LOD method
The LOD method is an efficient multiscale method developed on a coarse mesh $\mathcal{T}_{H}$ with a mesh size $\varepsilon\ll H$ and a quasi-local linear operator $I_{H}:V\to V_{H},\;I_{H}v=v$ for $\forall v\in V$. The computational finite element space $V_{H}$, with the local mesh size $H:=\max\limits_{T\in\mathcal{T}_{H}}\text{diam}(T)$ is defined by
$V_{H}:=\mathcal{H}^{1}_{\#}(\Omega)\cap\mathcal{Q}_{1}(\mathcal{T}_{H}).$
As we utilize quadrilateral meshes in this work, $\mathcal{Q}_{1}(\mathcal{T}_{H})$ is the space of piecewise polynomials of partial degree $\leq 1$. The fine space $V_{h}$ on which the reference solution is defined, is constructed similarly as $V_{H}$ but based on a refined mesh $\mathcal{T}_{h}$. We assume that the coarse mesh size $H$ is an integer multiple of the periodicity length $\varepsilon$, which together with the periodicity of the coefficient gives additional structure to our problem. In this study, we let the bounded, surjective projection operator $I_{H}$ to be constructed via
@@
where $\Pi_{H}:V\to\mathcal{Q}_{1}(\mathcal{T}_{H})$ is the local $L^{2}$-projection and $E_{H}$ is the averaging operator that maps the discontinuous functions in $\mathcal{Q}_{1}(\mathcal{T}_{H})$ to $V_{H}$ given by
@@
for any $v\in\mathcal{Q}_{1}(\mathcal{T}_{H})$ and vertex $z$ of $\mathcal{T}_{H}$. One more important ingredient in the LOD method is a domain localization based on element patches. For an element $T\in\mathcal{T}_{H}$ and $\mathfrak{k}\in\mathbb{N}_{0}$, the patch $U_{\mathfrak{k}}(T)\subset\Omega$ centered around $T$ is defined by
@@
The patch $U_{\mathfrak{k}}(T)$ is called the $\mathfrak{k}$-th layer patch. In the LOD method, we recover the fine-scale space on these local patches by the kernel of the operator $I_{H}$, i.e.
$W^{\mathfrak{k}}:=W_{H}(U_{\mathfrak{k}}(T)):=\{w\in W_{H}:\;\text{supp}(w)\subseteq U_{\mathfrak{k}}(T)\},$
where $W_{H}:=\text{ker}(I_{H}|_{V})=\{v\in V:I_{H}(v)=0\}\subset V$. Next, we define the element correction operator, $\mathscr{C}_{\mathfrak{k},T}(A):V\to W^{\mathfrak{k}}(T)$ that solves
$a(\mathscr{C}_{\mathfrak{k},T}v,w)=(A\nabla(\mathscr{C}_{\mathfrak{k},T}(A)v),\nabla w)_{U_{\mathfrak{k}}(T)}=(A\nabla v,\nabla w)_{T}=\int\limits_{T}(A\nabla v)\cdot\nabla wdx,\;\;\forall w\in W^{\mathfrak{k}}(T).$ (4)
The localized correction operator $\mathscr{C}_{\mathfrak{k}}(A):V\to W^{\mathfrak{k}}$ is the sum
$\mathscr{C}_{\mathfrak{k}}(A):=\sum\limits_{T\in\mathcal{T}_{H}}\mathscr{C}_{\mathfrak{k},T}.$ (5)
@@
The ideal correction operator $\mathscr{C}:V\to W_{H}$, i.e., the correction operator taken over the whole domain, is an $a$-orthogonal operator satisfying
$a((1-\mathscr{C})v,w)=0,\;\;\text{ for }v\in V\text{ and }w\in W_{H}.$
The localized multiscale space is then defined by
$V^{\mathrm{ms}}_{H,\mathfrak{k}}:=(1-\mathscr{C}_{\mathfrak{k}}(A))V_{H}=\text{span}\{\varphi_{z}-\mathscr{C}_{\mathfrak{k}}(A)\varphi_{z}\}_{z\in N}$ (6)
with the nodal basis $\{\varphi_{z}\}_{z\in N}$ of $V_{H}$. We can now introduce the LOD method.
The LOD method: Find $({\lambda}_{H,\mathfrak{k}},{u}_{H,\mathfrak{k}})\in\mathbb{R}\times V_{H}$ such that
${a}_{\mathfrak{k}}({u}_{H,\mathfrak{k}},(1-\mathscr{C}_{\mathfrak{k}})v)={\lambda}_{H,\mathfrak{k}}{m}_{\mathfrak{k}}({u}_{H,\mathfrak{k}},(1-\mathscr{C}_{\mathfrak{k}}){v})\;\;\forall{v}\in V_{H}$ (7)
@@
For further explanation on the LOD method, we refer to [MP14] and note the following results without proof.
@@
Let $I_{H}=E_{H}\circ\Pi_{H}$ be as defined in (3). Then, for $v\in V$ and $w\in W_{H}=\mathrm{kern}(I_{H})$,
@@
Without localization, i.e., $\mathfrak{k}=\infty$, the ideal LOD method for the problem (7), has been proposed and rigorously analyzed in [MP15, MP21]. Combining these estimates and proof techniques with Theorem 2.2, one arrives at the following error estimates for the eigenvalues and eigenfunctions of the LOD with localization
$\begin{split}\frac{{\lambda^{l}}_{H,\mathfrak{k}}-\lambda^{l}}{\lambda^{l}}&\lesssim\Big{(}g(\mathfrak{k})+H^{2}\Big{)}H^{2}\\ \norm{u^{(l+j)}-{u}_{H,\mathfrak{k}}^{(l+j),\mathrm{ms}}}_{L^{2}(\Omega)}&\lesssim\Big{(}g(\mathfrak{k})+H^{2}\Big{)}H\\ \norm{u^{(l+j)}-{u}_{H,\mathfrak{k}}^{(l+j),\mathrm{ms}}}_{\mathcal{A}}&\lesssim\Big{(}\sqrt{g(\mathfrak{k})}+H\Big{)}H.\end{split}$ (8)
By choosing $\mathfrak{k}$ large enough such that $g(\mathfrak{k})\sim H^{2}$, we obtain the same order of convergence in (8) as in [MP21].

### v2

@@ 1 Introduction
In this work, we aim to obtain accurate and efficient approximations of the solutions to the stochastic PDE for many different samples. However, in a multiscale setting, standard finite element methods are, in general, not feasible due to the required fine discretization. Therefore, computational multiscale methods are employed, where we focus on the Localized Orthogonal Decomposition (LOD) method [AHP21, MP21, HM14, EHMP19, MP14] in this work. The key idea is to construct problem-adapted basis functions incorporating information about the multiscale nature of the problem, so that a low-dimensional ansatz space yields sufficiently accurate solutions. However, solving the problem for many different samples remains quite expensive, as the problem-adapted basis functions must be re-computed for each realization of the (perturbed) coefficient.
The combined challenges of randomness and multiscale features have been tackled in several studies, often in the MC context and based on different multiscale methods. In [Owh17], the authors construct so-called gamblets – hierarchical basis functions that enable sparse and orthogonal multiresolution decomposition. The work in [ABS13] considers multilevel MC methods for problems with several, separated scales. In [BH17], a finite element-based hierarchical MC method is presented for problems with periodic multiscale coefficients generated by an ergodic dynamical system acting on a probability space. Multiscale model reduction is combined with data-driven approaches in [ZCH15] to obtain a reduced dimensional system in both physical and stochastic aspects. In [LBLT14], the authors propose a Multiscale Finite Element Method with a deterministic multiscale basis by taking into account the perturbative nature of the random coefficients, which have a similar structure as our model below. Very recently, [EKM24] suggests a combination of the LOD with neural networks for an efficient evaluation of the coefficient-to-basis map for general parametrized multiscale problems, that might be applicable in the random case as well. All the aforementioned works primarily consider the source problem. With regard to the eigenvalue problem, [XZO19] combines operator-adapted wavelets (gamblets) with hierarchical subspace correction to efficiently block-diagonalize the operator and refine solutions across scales. The algorithm computes eigenpairs through coarse-to-fine corrections in a multigrid fashion. A preconditioned inverse iteration with an optimal multigrid solver is used in [AP19] to efficiently compute approximations of the localized low-energy eigenstates to a linear random Schrödinger operator. Note that we cannot expect such localized eigenstates in our setting.
@@
$-\nabla\cdot(A(x,\omega)\nabla u(x,\omega))=\lambda(\omega)u(x,\omega)$ (1)
where $A(x,\omega)$ is a realization of the form
$A(x,\omega)=A_{\varepsilon}(x)+b_{p}(x;\omega)B_{\varepsilon}(x).$ (2)
Here, $A_{\varepsilon}(x)$ and $B_{\varepsilon}(x)$ are deterministic, multiscale coefficients, and $b_{p}(x;\omega)$ follows a Bernoulli law with probability $p$. Detailed assumptions are postponed to Section 2.1. The random defects model for $A$ is the same as in [MV22] and originally considered in [ALB11, ALB12]. Figure 1 represents two example coefficients of $A$ which fall into the above model. In the random checkerboard (Figure 1(a)), the coefficient $A$ is generated by randomly including $\varepsilon$-length square-spots, i.e. defects, with a probability $p$ on to the constant background coefficient $A_{\varepsilon}$. In random erasure (Figure 1(b)), $A_{\varepsilon}$ is made of square-spots that are periodically arranged with periodicity $\varepsilon$ on a constant background value. Then the coefficient $A$ is generated by randomly removing the $\varepsilon$-length square-spots with a probability $p$. While in the random checkerboard a defect is represented by the random inclusion of a square-spot, in the random erasure a defect is indicated by the random removal of a square-spot.
@@ 2.1 Model problem
In this section, we describe the problem for a single sample $\omega$. For simplicity, we thus omit the random variable $\omega$ in the following. Recall, however, that the final goal is to find an effective way of solving this problem for many samples. We assume the domain $\Omega\subset\mathbb{R}^{d}$ to be a unit hypercube, i.e., $\Omega=[0,1]^{d}$ with $d\in\{1,2,3\}$ where we impose periodic boundary conditions. When we later discretize our problem, we always use Cartesian meshes.
The weak form of the elliptic eigenvalue problem in (1) is given by: Find $(\lambda,u)\in\mathbb{R}\times V$ such that
@@
Note that although equation (3) is stated in a general form for simplicity, it encloses all eigenpairs $(\lambda^{(l+j)},u^{(l+j)})$ for $j=0,\cdots,r-1$, accounting for multiplicities of the $l$-th eigenvalue. This interpretation applies similarly to all subsequent discretized formulations. The functional space $V$ is the conforming finite element space given by
$V:=V_{h}\subseteq\mathcal{H}^{1}_{\#}(\Omega):=\left\{v\in H^{1}_{\text{loc}}\left(\mathbb{R}^{d}\right)\ \middle|\ v\text{ is }\Omega\text{-periodic and }v|_{\Omega}\in H^{1}(\Omega)\right\},$
where $H^{1}(\Omega):=\{v\in L^{2}(\Omega):\nabla v\in(L^{2}(\Omega))^{d}\}$ is the usual Sobolev space, and $H^{1}_{\text{loc}}(\mathbb{R}^{d})$ consists of functions in $H^{1}(\mathcal{U})$ for every bounded open set $\mathcal{U}\subset\mathbb{R}^{d}$. The $\Omega$-periodicity means that the function $v$ can be extended to a function on $\mathbb{R}^{d}$ that is periodic with respect to the domain $\Omega$. We assume the mesh size $h$ to be small enough that we treat the solution $u\in V_{h}$ as the reference solution, i.e., the error towards the exact solution in #$\mathcal{H}^{1}_{\#}(\Omega)$ is negligible. The energy norm and the $L^{2}$-norm are as usually defined by $a(\cdot,\cdot)=:{\norm{\cdot}^{2}_{\mathcal{A}}}={\norm{A^{1/2}\nabla\cdot}^{2}_{L^{2}(\Omega)}}$ and $m(\cdot,\cdot)=:\norm{\cdot}^{2}_{L^{2}(\Omega)}$ respectively. The realization $A(x)\in L^{\infty}(\Omega,\mathbb{R})$ is a sample of a randomly perturbed coefficient which is uniformly bounded and elliptic
@@
For simplicity, we assume in the following that $A(x)$ is scalar-valued; however, all results can be immediately transferred to the matrix-valued case. The underlying deterministic coefficients $A_{\varepsilon}$ and $B_{\varepsilon}$ of $A(x)$ given in (2) are taken to be spectrally bounded as
@@
$0<\alpha\leq\text{ess}\inf\limits_{x\in\Omega}(A_{\text{per}}(x)+B_{\text{per}}(x)),\;\;\;\;\infty>\beta\geq\text{ess}\sup\limits_{x\in\Omega}(A_{\text{per}}(x)+B_{\text{per}}(x))$
@@
Here, the function $\chi$ is the characteristic function of each defect with $Q\subseteq[0,1]^{d}$ and $I:=\{\kappa\in\mathbb{Z}^{d}\ |\varepsilon(\kappa+Q)\subset\Omega\}$. The independent random variables $\hat{b}^{i}_{p}$ are Bernoulli distributed with probability $p$. Precisely, $\hat{b}^{i}_{p}=0$ with probability $1-p$ and $\hat{b}^{i}_{p}=1$ with probability $p$.
@@
For conciseness, we use the notation $a\lesssim b$ to denote the relationship $a\leq Cb$ where the generic constant $C$ is independent of the discretization parameters $H,h,\mathfrak{k}$ as well as $\varepsilon$ throughout the rest of the paper.
@@ 2.2 The LOD method
The LOD method is a numerical approach designed for solving elliptic differential equations with highly heterogeneous or multiscale coefficients. It enables us to accurately approximate solutions without requiring a fine mesh over the whole domain. Instead, it works on a coarse mesh while incorporating localized fine-scale corrections. In this section, let us briefly outline the LOD method.
We start by introducing a coarse mesh $\mathcal{T}_{H}$ over the domain $\Omega$, with a local mesh size $H:=\max\limits_{T\in\mathcal{T}_{H}}\text{diam}(T)$. The fine-scale behavior is assumed to occur on a smaller scale $\varepsilon\ll H$. We define the computational finite element space $V_{H}$, consisting of continuous, piecewise bilinear functions on quadrilateral (or cuboidal) elements by
$V_{H}:=\mathcal{H}^{1}_{\#}(\Omega)\cap\mathcal{Q}_{1}(\mathcal{T}_{H})$
where $\mathcal{Q}_{1}(\mathcal{T}_{H})$ is the space of piecewise polynomials of partial degree $\leq 1$.
The fine space $V_{h}$ on which the reference solution is defined, is constructed similarly as $V_{H}$ but based on a refined mesh $\mathcal{T}_{h}$. We assume that the coarse mesh size $H$ is an integer multiple of the periodicity length $\varepsilon$, which together with the periodicity of the coefficient gives additional structure to our problem.
To distinguish coarse- and fine-scale features, the LOD method utilizes a quasi-local linear operator $I_{H}:V\to V_{H}$. The kernel space
$W:=\text{ker}(I_{H}|_{V})=\{v\in V:I_{H}(v)=0\}\subset V$
captures the fine-scale components unresolved by $V_{H}$.
In this study, we define this bounded, surjective projection operator $I_{H}$ as follows:
@@
where $\Pi_{H}:V\to\mathcal{Q}_{1}(\mathcal{T}_{H})$ denotes the local $L^{2}$-projection onto the piecewise linear discontinuous finite element space, and $E_{H}$ is the averaging operator that maps discontinuous functions in $\mathcal{Q}_{1}(\mathcal{T}_{H})$ to the conforming space $V_{H}$ defined by
@@
for any $v\in\mathcal{Q}_{1}(\mathcal{T}_{H})$ and vertex $z$ of $\mathcal{T}_{H}$. This interpolation operator $I_{H}=E_{H}\circ\Pi_{H}$ is $L^{2}$-stable as discussed in [PS17], that is,
$\norm{I_{H}v}_{L^{2}(\Omega)}\leq C_{\text{stab}}\norm{v}_{L^{2}(\Omega)}\text{ for all }v\in V$
with a constant $C_{\text{stab}}$ that depends only on the mesh regularity.
Having established the construction and stability of the interpolation operator $I_{H}$, we now turn to the construction of the LOD method. In the ideal setting, fine-scale corrections are computed globally by a corrector operator $\mathscr{C}(A):V\to W$ defined over the entire domain. This operators is defined via its $a$-orthogonal property
$a((1-\mathscr{C}(A))v,w)=0,\;\;\text{ for }v\in V\text{ and }w\in W.$
The key idea is to use the finite-dimensional space $(1-\mathscr{C}(A))V_{H}$ in a Galerkin scheme for (3). Although computing global correctors is computationally expensive, their exponential decay from the associated coarse elements enables efficient local approximation within small patches of the domain. We introduce these patches as follows.
For an element $T\in\mathcal{T}_{H}$ and $\mathfrak{k}\in\mathbb{N}_{0}$, the patch $U_{\mathfrak{k}}(T)\subset\Omega$ centered around $T$ is defined by
@@
The patch $U_{\mathfrak{k}}(T)$ is called the $\mathfrak{k}$-th layer patch. The localized fine-scale space associated with a patch $U_{\mathfrak{k}}(T)$ is then defined as
$W^{\mathfrak{k}}:=W(U_{\mathfrak{k}}(T)):=\{w\in W:\;\text{supp}(w)\subseteq U_{\mathfrak{k}}(T)\}.$
The localized corrector problem defined on $W^{\mathfrak{k}}$ is designed to compute a fine-scale correction that removes unresolved oscillations from any coarse function. In the following, we define the element correction operator, $\mathscr{C}_{\mathfrak{k},T}(A):V\to W^{\mathfrak{k}}$ that solves
$a\left(\mathscr{C}_{\mathfrak{k},T}(A)v,w\right)=\left(A\nabla(\mathscr{C}_{\mathfrak{k},T}(A)v),\nabla w\right)_{U_{\mathfrak{k}}(T)}=\left(A\nabla v,\nabla w\right)_{T}=\int\limits_{T}(A\nabla v)\cdot\nabla wdx,\;\;\forall w\in W^{\mathfrak{k}}.$
Observe that the space $W^{\mathfrak{k}}$ is constructed by restricting the global fine-scale space $W$ to the localized patches where $I_{H}$ is our continuous quasi-interpolation operator. Since the kernel of a continuous linear operator on a Hilbert space is closed, and the restriction preserves closedness, it follows that $W^{\mathfrak{k}}\subset\mathcal{H}^{1}_{\#}(\Omega)$ is also closed. Consequently, the element correction operator defined through a local linear variational problem on $W^{\mathfrak{k}}$ is well-defined by the Lax-Milgram theorem.
Next we define the localized correction operator $\mathscr{C}_{\mathfrak{k}}(A):V\to W^{\mathfrak{k}}$ as the sum
$\mathscr{C}_{\mathfrak{k}}(A):=\sum\limits_{T\in\mathcal{T}_{H}}\mathscr{C}_{\mathfrak{k},T}(A).$
@@
We stress stress that the localized correction operator $\mathscr{C}_{\mathfrak{k}}$ does not inherit the $a$-orthogonality property of the ideal correction operator $\mathscr{C}(A)$ due to the truncation errors. In addition to being $a$-orthogonal, the ideal corrector $\mathscr{C}(A)$ is also stable in the energy norm for all $v\in V$, as it acts as an $a$-orthogonal projection onto the fine-scale space $W$. Since the orthogonal projections in Hilbert spaces do not increase the norm, $\norm{\mathscr{C}(A)v}_{\mathcal{A}}$ is guaranteed to be bounded by $\norm{v}_{\mathcal{A}}$. Although the localized corrector $\mathscr{C}_{\mathfrak{k}}(A)$ lacks the $a$-orthogonality, it is stable in the energy norm for any $v\in V_{H}$. This follows because the element correction operator is well defined and $\mathscr{C}_{\mathfrak{k},T}(A)v$ solves a coercive problem on a patch where these patches have uniformly bounded overlap.
We now define the localized multiscale space by
$V^{\mathrm{ms}}_{H,\mathfrak{k}}:=\left(1-\mathscr{C}_{\mathfrak{k}}(A)\right)V_{H}=\text{span}\{\varphi_{z}-\mathscr{C}_{\mathfrak{k}}(A)\varphi_{z}\}_{z\in N}$ (5)
where $N$ denotes the set of vertices of $\mathcal{T}_{H}$ and $\{\varphi_{z}\}_{z\in N}$ is the nodal basis of $V_{H}$.
From this point onward, we abbreviate $\mathscr{C}(A)=\mathscr{C},\mathscr{C}_{\mathfrak{k}}(A)=\mathscr{C}_{\mathfrak{k}}$ and $\mathscr{C}_{\mathfrak{k},T}(A)=\mathscr{C}_{\mathfrak{k},T}$. We can now introduce the LOD problem.
The LOD problem: Find $({\lambda}_{H,\mathfrak{k}},{u}_{H,\mathfrak{k}})\in\mathbb{R}\times V_{H}$ such that
${a}_{\mathfrak{k}}({u}_{H,\mathfrak{k}},v)={\lambda}_{H,\mathfrak{k}}{m}_{\mathfrak{k}}({u}_{H,\mathfrak{k}},v)\;\;\forall{v}\in V_{H}$ (6)
@@
For a comprehensive treatment of the LOD method, we refer the reader to [MP21], and include the following results [MP21, Theorem 4.3, Lemma 8.1] without proof.
@@
Let $I_{H}=E_{H}\circ\Pi_{H}$ be as defined in (4). Then, for $v\in V$ and $w\in W=\mathrm{kern}(I_{H})$,
@@
Without localization, i.e., $\mathfrak{k}=\infty$, the ideal LOD method for the problem (6), has been proposed and rigorously analyzed in [MP15, MP21]. Combining these estimates and proof techniques with Theorem 2.2, one arrives at the following error estimates for the $l$-th eigenvalue solution of the LOD with localization
$\begin{split}\frac{\lambda^{(l)}_{H,\mathfrak{k}}-\lambda^{(l)}}{\lambda^{l}}&\lesssim\left(g(\mathfrak{k})+H^{2}\right)H^{2}\\ \norm{u^{(l)}-{u}_{H,\mathfrak{k}}^{(l),\mathrm{ms}}}_{L^{2}(\Omega)}&\lesssim\left(g(\mathfrak{k})+H^{2}\right)H\\ \norm{u^{(l)}-{u}_{H,\mathfrak{k}}^{(l),\mathrm{ms}}}_{\mathcal{A}}&\lesssim\left(\sqrt{g(\mathfrak{k})}+H\right)H\end{split}$ (7)
where $u^{(l),\text{ms}}_{H,\mathfrak{k}}=(1-\mathscr{C}_{\mathfrak{k}})u^{(l)}_{H,\mathfrak{k}}$. By choosing $\mathfrak{k}$ large enough such that $g(\mathfrak{k})\approx H^{2}$, we obtain the same order of convergence in (7) as established in [MP21, Theorems 8.4, 8.5].